# Entity Alignment with Noisy Annotations from Large Language Models

### v1

@@ Abstract
Entity alignment (EA) aims to merge two knowledge graphs (KGs) by identifying equivalent entity pairs. While existing methods heavily rely on human-generated labels, it is prohibitively expensive to incorporate cross-domain experts for annotation in real-world scenarios. The advent of Large Language Models (LLMs) presents new avenues for automating EA with annotations, inspired by their comprehensive capability to process semantic information. However, it is nontrivial to directly apply LLMs for EA since the annotation space in real-world KGs is large. LLMs could also generate noisy labels that may mislead the alignment. To this end, we propose a unified framework, LLM4EA, to effectively leverage LLMs for EA. Specifically, we design a novel active learning policy to significantly reduce the annotation space by prioritizing the most valuable entities based on the entire inter-KG and intra-KG structure. Moreover, we introduce an unsupervised label refiner to continuously enhance label accuracy through in-depth probabilistic reasoning. We iteratively optimize the policy based on the feedback from a base EA model. Extensive experiments demonstrate the advantages of LLM4EA on four benchmark datasets in terms of effectiveness, robustness, and efficiency.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have showcased their superior capability in processing semantic information, which has significantly advanced various graph learning tasks such as node classification (Chen et al., 2024b), graph reasoning (Zhao et al., 2023a; Chai et al., 2023), recommender systems (Zhou et al., 2022; Wu et al., 2023), SQL query generation (Zhang et al., 2024a), and knowledge graph-based question answering (Wang et al., 2024; Zhang et al., 2024b; Dong et al., 2024). Their capacity to extract meaningful insights from graph data opens up new possibilities for automating EA. Notably, recent studies (Zhong et al., 2022; Zhao et al., 2023b; Jiang et al., 2024) have explored the use of LLMs in EA, primarily focusing on finetuning a pretrained LLM such as Bert to learn semantic-aware representations, relying on accurate seed alignments as training labels. Yet, the potential of LLMs for label-free EA via in-context learning remains unexplored.
@@ 4 Experiments
In this section, we conduct experiments to evaluate the effectiveness of our framework. We begin by introducing the experimental settings. Then, we present experiments to answer the following research questions: RQ1. How effective is the overall framework? RQ2. What is the impact of the choice of LLM on the cost and performance of LLM4EA? RQ3. What is the effect of the label refiner? RQ4. What is the impact of active selection?
@@ 4.2.3 Ablation study
The runtime overhead is directly proportional to the number of active selection iterations, $n$, since each iteration involves a subsequent label refinement process. To explore the runtime-performance trade-off in the LLM4EA approach, we examine the Pareto frontier of runtime versus performance. We conduct entity alignment experiments with a fixed query budget, varying the number of active selection iterations. The results of these experiments are illustrated in Figure 4.
@@ 6 Limitations
Firstly, during active selection, we distribute the query budget evenly for each selection rather than dynamically customizing the budget allocation for each iteration. This allocation approach could be improved by developing a more adaptive budget allocation strategy. Secondly, the framework currently does not provide direct support for temporal KGs. Although the probabilistic reasoning and active selection components inherently support entity alignment on evolving KGs, the base EA model is transductive. This necessitates retraining the model whenever new entities and relation triples are introduced into the KGs. However, this can be complemented by the research line of inductive entity alignment, such as path-based embedding models or logic-based models, which can generalize to previously unseen entities without the need for retraining.

### v2

@@ Abstract
Entity alignment (EA) aims to merge two knowledge graphs (KGs) by identifying equivalent entity pairs. While existing methods heavily rely on human-generated labels, it is prohibitively expensive to incorporate cross-domain experts for annotation in real-world scenarios. The advent of Large Language Models (LLMs) presents new avenues for automating EA with annotations, inspired by their comprehensive capability to process semantic information. However, it is nontrivial to directly apply LLMs for EA since the annotation space in real-world KGs is large. LLMs could also generate noisy labels that may mislead the alignment. To this end, we propose a unified framework, LLM4EA, to effectively leverage LLMs for EA. Specifically, we design a novel active learning policy to significantly reduce the annotation space by prioritizing the most valuable entities based on the entire inter-KG and intra-KG structure. Moreover, we introduce an unsupervised label refiner to continuously enhance label accuracy through in-depth probabilistic reasoning. We iteratively optimize the policy based on the feedback from a base EA model. Extensive experiments demonstrate the advantages of LLM4EA on four benchmark datasets in terms of effectiveness, robustness, and efficiency. Codes are available via https://github.com/chensyCN/llm4ea_official.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have showcased their superior capability in processing semantic information, which has significantly advanced various graph learning tasks (Li et al., 2023) such as node classification (Chen et al., 2024b; Li et al., 2024), graph reasoning (Zhao et al., 2023a; Chai et al., 2023), recommender systems (Zhou et al., 2022; Wu et al., 2023), SQL query generation (Zhang et al., 2024a), and knowledge graph-based question answering (Wang et al., 2024; Zhang et al., 2024b; Dong et al., 2024). Their capacity to extract meaningful insights from graph data opens up new possibilities for automating EA. Notably, recent studies (Tang et al., 2020; Zhong et al., 2022; Zhao et al., 2023b; Jiang et al., 2024) have explored the use of LLMs in EA, primarily focusing on finetuning a pretrained LLM such as Bert to learn semantic-aware representations, relying on accurate seed alignments as training labels. Yet, the potential of LLMs for label-free EA via in-context learning remains unexplored.
@@ 4 Experiments
In this section, we conduct experiments to evaluate the effectiveness of our framework. We begin by introducing the experimental settings. Then, we present experiments to answer the following research questions: RQ1. How effective is the overall framework? RQ2. What is the impact of the choice of LLM on the cost and performance of LLM4EA? RQ3. What is the effect of the label refiner? RQ4. What is the impact of active selection? RQ5. How is the runtime-performance trade-off managed?
@@ 4.2.3 Ablation study
We answer RQ5. in this experiment. The runtime overhead is directly proportional to the number of active selection iterations $n$, since each iteration involves a subsequent label refinement process. To explore the runtime-performance trade-off, we examine the Pareto frontier of runtime versus performance. We conduct entity alignment experiments with a fixed query budget, varying the number of active selection iterations. The results of these experiments are illustrated in Figure 4.
@@ 6 Limitations
Firstly, during active selection, we distribute the query budget evenly for each selection rather than dynamically customizing the budget allocation for each iteration. This allocation approach could be improved by developing a more adaptive budget allocation strategy. Secondly, the framework currently does not provide direct support for temporal KGs. Although the probabilistic reasoning and active selection components inherently support entity alignment on evolving KGs, the base EA model is transductive. This necessitates retraining the model whenever new entities and relation triples are introduced into the KGs. However, this limitation can be complemented by the research line of inductive entity alignment, such as path-based embedding models or logic-based models, which can generalize to previously unseen entities without the need for retraining.